OpenStudy (anonymous):

f(x)=x^3+6x^2+x^1/2 and g(x)=x^1/2 find f(x)/g(x)

OpenStudy (xapproachesinfinity):

well what's stopping you do f(x)/g(x)!

OpenStudy (jdoe0001):

\(\bf \large { f(x)=x^3+6x^2+x^{\frac{1}{2}} \qquad g(x)=x^{\frac{1}{2}} \\ \quad \\ \cfrac{f(x)}{g(x)}\implies \cfrac{x^3+6x^2+x^{\frac{1}{2}}}{x^{\frac{1}{2}}}\implies \cfrac{x^3}{x^{\frac{1}{2}}}+\cfrac{6x^2}{x^{\frac{1}{2}}}+\cfrac{x^{\frac{1}{2}}}{x^{\frac{1}{2}}} \\ \quad \\ recall\implies {\color{brown}{ \cfrac{a^m}{a^n}\implies a^{n-m}}} }\)

OpenStudy (jdoe0001):

hmmm lemem fix that a bit \(\bf \large { f(x)=x^3+6x^2+x^{\frac{1}{2}} \qquad g(x)=x^{\frac{1}{2}} \\ \quad \\ \cfrac{f(x)}{g(x)}\implies \cfrac{x^3+6x^2+x^{\frac{1}{2}}}{x^{\frac{1}{2}}}\implies \cfrac{x^3}{x^{\frac{1}{2}}}+\cfrac{6x^2}{x^{\frac{1}{2}}}+\cfrac{x^{\frac{1}{2}}}{x^{\frac{1}{2}}} \\ \quad \\ recall\implies {\color{brown}{ \cfrac{a^m}{a^n}\implies a^{m-n}}} }\)